# Reconciling Inflation with Hubble Anisotropies

### v1

@@ 9. Conclusion
We have seen however that, if we introduce torsion into cosmology, then there is a price to be paid : in all but a few special cases, Inflation no longer completely isotropises the early Universe. But recent observations show that this may be precisely what we need —  if indeed one or more of the various kinds of cosmic anisotropy should be confirmed observationally, then torsion may be a way to reconcile such observations with Inflation. Not only that : if this works, torsion will clearly be by far the most natural way of effecting that reconciliation.
However, the operative word here is, as it should be, “if ”. We have seen that torsion can, extremely naturally, account for some forms of anisotropy, but by no means all. Anisotropy along more than one axis ; anisotropy that involves no rotation ; anisotropy such that the magnitude of the Hubble field is not minimal along the distinguished axis ; anisotropy not accompanied by anomalies in the early value of the Hubble parameter : any of these, if confirmed, might present a serious challenge to our proposed reconciliation of Inflation with anisotropy.

### v2

@@ 9. Conclusion
We have seen however that, if we introduce torsion into cosmology, then there is a price to be paid : in all but a few special cases, Inflation no longer completely isotropises the early Universe. But recent observations show that this may be precisely what we need: the evidence for anisotropy is hard to ignore [24]. If one or more of the various kinds of cosmic anisotropy should be confirmed observationally, then torsion may be a way —  or the way —  to reconcile such observations with Inflation.
However, the operative word here is, as it should be, “may ”. We have seen that torsion can, extremely naturally, account for some forms of anisotropy, but by no means all. Anisotropy (at the end of Inflation) along more than one axis, or along one axis but not minimal there ; anisotropy that involves no rotation, or is such that the amount of anisotropy is not correlated with the amount of rotation ; anisotropy not accompanied by anomalies in the early value of the Hubble parameter : any of these, if confirmed, might present a serious challenge to our proposed reconciliation of Inflation with anisotropy.